HUBBLE SPACE TELESCOPE DAILY REPORT #3045 PERIOD COVERED: 0000Z (UTC) 01/29/02 - 0000Z (UTC) 01/30/02 Daily Status Report as of 030/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD/MA2 9161 (The Ultraviolet Properties of Obscured QSOs) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to observe a subset of The Two Micron All-Sky Survey {2MASS} -discovered QSOs to: 1} determine if UV absorption- line features are present as seen in several other highly polarized AGN populations; 2} test relationships between X-ray and UV/optical attenuation, AGN spectral type, optical polarization, and detected UV absorption-line features; and 3} characterize the spectral energy distributions and UV properties of obscured QSOs. There were no problems. 1.2 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.3 Completed Four Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.4 Completed STIS/MA1 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.5 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.6 Completed WF/PC-2 9118 (Tracing the Cosmic Expansion to z>1 with Type Ia Supernovae) The WF/PC-2 was used to observe type Ia supernovae in order to provide evidence for an accelerating universe. The case for cosmic acceleration rests almost entirely on the observation that the observed SN Ia at z~0.5 are 0.25 magnitudes fainter than expected for a non-accelerating Universe. It is proposed to follow five SN Ia in the range 0.95 0 cosmology, this experiment is a powerful and straightforward way to assess the reliability of the SN Ia measurements. In addition, if SN Ia are reliable standard candles, the proposed observations will significantly increase the precision with which Omega_Lambda and Omega_M are measured. There was no reported anomalous activity. 1.7 Completed STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.8 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.9 Completed STIS/CCD 9239 (Absolute Proper Motions of Nearby Dwarf Spheroidal Galaxies-Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to measure precise absolute proper motions for four dwarf spheroidal satellites of the Milky Way using spectroscopically-confirmed background QSOs to define a zero- velocity reference frame. The proposal completed with no reported problems. 1.10 Completed Two Sets of STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space Telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.11 Completed WF/PC-2 8059 (POMS Test Proposal: Targeted Parallel Archive Proposal) The WF/PC-2 was used to observe the parallel opportunities available in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local Group galaxies offer the opportunity for a closer look at young stellar populations. Narrow-band images in F656N can be used both to identify young stars via their emission lines, and to map the gas distribution in star-forming regions. The observations completed nominally. 1.12 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.13 Completed STIS/CCD 9187 (A Lyman Limit Snapshot Survey: The Search for Candidate D/H Absorbers) The Space Telescope Imaging Spectrograph (CCD) was used to measure D/H in gas clouds seen against background QSOs. This provides a unique and fundamental cosmological probe of the baryonic density parameter. To date, most of the QSO absorption line D/H constraints have been derived at high redshift. In this proposal the proposers investigate a lower redshift range, because there are two potentially crucial advantages to be gained by carrying out D/H measurements at intermediate redshift. First, the background of HI forest lines is lower and contamination of the DI feature is significantly reduced. Second, a far larger sample of very bright QSOs exists. There were no reported problems. 1.14 Completed STIS/CCD/MA1 9137 (Quasar Absorbers and Large Scale Structure) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to perform spectroscopy of 15 bright quasars in a 22 square degree region that has well-sampled galaxy redshifts. No problems were encountered. 1.15 Completed WF/PC-2 9319 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal, designed to simulate future scientific plans. The proposal completed without incident. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: The SSR-1 EDAC error counter was cleared twice per ROP SR-1A. Using ROP NS-3, the NSSC-1 status buffer was dumped at 029/2000Z. A TTR was generated when there was a required re-transmit at 030/0746Z during a 486 uplink. The engineering status buffer limit was updated (SESBSLD=1) at 030/0752Z in accordance with ROP DF-18A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW